OpenStudy (anonymous):

Can someone please help me understand this question? If y varies directly with x and y = -7 when x = -1, find x when y = -84. A. 84 B. 7 C.-12 I got A (84)

OpenStudy (anonymous):

Direct variation means y=kx Set up a proportion. \[\frac{ -7 }{ -1 }=\frac{ -84 }{ x }\] -84 x -1= 84 84 / -7= -12 so it's C

OpenStudy (anonymous):

oh, okay thank you so much :)